OpenStudy (anonymous):

through (-3-1) parallel to y=x-4

OpenStudy (kkutie7):

parallel means same slope ok

OpenStudy (kkutie7):

so the slope of y=x-4 is 1

OpenStudy (kkutie7):

now you use (-3,-1) y=slope(x-x1)+y y=1x-(-3)-1 y=x+3-1 y=x+2..... check my work to be sure ok =)